Problem 1. The sum of the digits of a two-digit number is 14. If the digits are reversed, the number is increased by 18. What is the number?
....
Problem 2. One number is less than another number. The sum of the two numbers is four times their difference. What are the numbers?

Answer by stanbon(75887)

Problem 1. The sum of the digits of a two-digit number is 14. If the digits are reversed, the number is increased by 18. What is the number?
------
Let the number be 10t+u
The reversed number is 10u+t
Equations:
t + u = 14
10u+t - (10t+u) = 18
---------------------------
Simplify the 2nd equation
t + u = 14
u - t = 2
---
Add to get 2u = 16
u = 8
Then t = 6
---
1st number: 68
....
Problem 2. One number is less than another number. The sum of the two numbers is four times their difference. What are the numbers?
---
Not enough information.
=============================
Cheers,
Stan H.